OpenStudy (anonymous):

The length of a rectangle is 3 inches less than twice the width. If the perimeter of the rectangle is 18 inches, find the width of the rectangle. __Inches

OpenStudy (anonymous):

Let: Breadth of rectangle=b According to the question, Length is 3 inches less than twice the width Therefore, Length of the rectangle=2b-3 Perimeter of a rectangle=2(Length)+2(Breadth)=2(Length+Breadth) But given perimeter is 18 Hence, 2[(2b-3)+b]=18 Solving this equation we get, Length=5 inches and Breadth=4 inches